40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1969     Entry time: Mon Sep 7 23:18:01 2009
Author: Alberto 
Type: Update 
Category: Locking 
Subject: Some locking attempts 

Tried to lock the interferometer but arm power didn't get over 65.

Tonight, after the weekend, I resumed the work on locking.

When I started the Mode Cleaner was unlocked because the MZ was also unlocked.

I aligned the MZ and the transmitted power reached about 2.5

Initially the interferometer lost lock at arm power of about 3-4. It looked like the alignment wasn't good enough. So I ran the alignment scripts a few times, first the scripts for the single parts and in the end the one for the full IFO.

Then I also locked again the MZ and this time the transmitted power got to about 4.

In the following locking attempts the the arm power reached 65 but then the lock got lost during the handing of CARM to C1:LSC-PD11_I

I'll keep working on that tomorrow night.

ELOG V3.1.3-